solve 6=e^0.2t. round your answer to the nearest hundredth
Answer by Alan3354(21605)

solve 6=e^0.2t. round your answer to the nearest hundredth
-------------------
0.2t = ln(6)
t = 5*ln(6)
t =~ 8.96